There were 415 tickets purchased for a major league baseball game. The upper level  tickets cost ​$12.50 and the tickets cost ​$8.00 The total amount of money spent was ​$4085.00 How many of each kind of ticket were​ purchased?
Answer by ikleyn(53937)

.
x tickets        at the price $12.50, and

(415-x) tickets  at the price  $8.00.


Write the total cost equation

    12.50x + 8*(415-x) = 4085  dollars.


Simplify and find x

    12.50x + 3320 - 8x = 4085,

    12.50x - 8x = 4085 - 3320

         4.50x  =    765

             x  =    765/4.50 = 170.


ANSWER.  170 tickets at the price of $12.50  and the rest, 415-170 = 245 tickets at the price of $8.


CHECK.  12.50*170 + 8*245 = 4085,  total cost.  ! correct !

Solved.

-------------------

To see many other similar solved problems using one equation in one unknown,  look into the lesson
    - Typical problems on buying and selling items
in this site.

To see many other similar solved problems using systems of two equations in two unknowns,  look into the lesson
    - Using systems of equations to solve problems on tickets
in this site.